Young’s modulus of a wire is Y, strain energy per unit volume is E then its strain is given by what?

Grade:11

2 Answers

Arun
3 years ago

as energy stored=1/2*stress*strain*volume

and energy/volume=E (given)

so, E=strain²*Y (strain=stress/Y ,Y=young mod. of elasticity)

so strain=(2E/Y)½

Vikas TU
3 years ago
Dear student 
as energy stored=1/2*stress*strain*volume
and energy/volume=E (given)
so, E=strain²*Y (strain=stress/Y ,Y=young mod. of elasticity)
so strain=(2E/Y)½
Good Luck